# Leakage and Interpretability in Concept-Based Models

### v1

@@ A.1 Datasets
For additional details on these datasets, we refer the reader to .
@@
This dataset is constructed by sampling a 3-dimensional latent variable $z\sim\mathcal{N}(0,\Sigma(\delta))$, with correlation matrix $\Sigma(\delta)_{ij}=\delta_{ij}+\delta(1-\delta_{ij})$ with $i,j=1,2,3$, $\delta_{ij}$ Kronecker symbol and $\delta\in(-1,1)$ indicating the correlation between components. The 7-dimensional input $x$ is a trigonometric function of the latent variables (see Mahinpei et al., 2021), while the three binary concepts are defined as the sign of the latent variables, $c_{i}=(z_{i}>0)$. The binary task label is the following linear function of concepts,
@@
For our experiments we generate a dataset of size 10K, with a 0.7/0.2/0.1 train / validation / test ratio. For Figures 2 and 3 we consider a simplified version of TabularToy($\delta$) with only the first two concepts $c_{1}$ and $c_{2}$ for each input, and binary task ${y=(c_{1}+c_{2}\geq 1)}$.
@@
In TabularToy(0.25), we remove $c_{3}$ from the three concepts. In dSprites(0), we remove $c_{4}$ and $c_{5}$ from the five concepts. In 3dshapes(0), we remove $c_{3}$ and $c_{6}$ from the six concepts.
@@
We provide here more details on the experiments to assess the effects of a misspecified final head discussed in Section 7. To generate non-linear concepts-task dependences in TabularToy($\delta$), we add a non-linear term to the original linear task in (13), resulting in the binary classification problem,
@@
The original tasks in (14) and (15) for dSprites($\gamma$) and 3dshapes($\gamma$) are non-linear, however a linear classifier is able to reach essentially perfect task accuracy (Table 3). To induce a sizeable misspecification, we thus add further higher-order non-linear terms. The tasks implemented for our experiments on dSprites(0) and 3dshapes(0) read,
@@ A.2 Model architectures and training
On TabularToy($\delta$) we use as concept encoder a 4-layer leaky-ReLU MLP with activations $\{7,64,64,3\}$, and a linear classifier head. On dSprites($\gamma$) and 3dshapes($\gamma$) we use a ResNet-18 encoder, and a 4-layer ReLU MLP with activations $\{k,64,64,\ell\}$. We train CEMs with 16-dimensional embeddings.
We train all models for 200 epochs, using the Adam optimiser with learning rate $10^{-3}$ and momentum 0.9. For hard models, we train the encoder and the classifier head for 200 and 20 epochs respectively. We train with batch sizes 512 for TabularToy($\delta$) and 32 for dSprites($\gamma$) and 3dshapes($\gamma$).
@@
A.3 Scores evaluation
Except for Figure 14, all the displayed scores are evaluated on test sets. The reported intervention performances are obtained with a random policy. In Figures 3, 4 and 16 we show results for single models, and the displayed means and 95% confidence intervals refer to a 5-fold evaluation of each score. In the rest of the paper, we repeat the evaluation of each score 5 times for individual models, and we display means and 95% confidence intervals representing the distribution of the mean upon 5-fold training for each model class. MIs and entropies are estimated via the KSG estimator (Kraskov et al., 2004) with 3 nearest neighbours. The OIS and NIS are evaluated using their original implementations and default settings from Espinosa Zarlenga et al. (2023a).
Appendix B Undesired features of the Niche Impurity Score
Being computed as an AUC, NIS = 1/2 indicates that concept representation subsets are not predictive of the value of concepts, while NIS = 1 is meant to indicate that the higher-order interconcept leakage is maximal. For a fair comparison of our scores with NIS, we plot $\widetilde{\text{NIS}}=2(\text{NIS}-1/2)$, taking values in $(0,1)$ corresponding to no and maximal leakage respectively.
Evaluating NIS on the pairs of models considered in Figure 4, this score appears to be anticorrelated with intervention performance and leakage (Figure 16). Furthermore, it takes proportionally very high values regardless of the model and dataset. This is confirmed by its behaviour on hard CBMs (Figure 15): the NIS is non-vanishing and very high in all cases. These undesired features that NIS generally exhibits signal potential flaws in its design and make it an unsuitable metric of leakage.
@@ Appendix B Undesired features of the Niche Impurity Score
Figure 16: $\widetilde{\text{NIS}}$ and leakage scores for the pairs of models assessed in Figure 4, Table 1.
@@ Appendix C Concept-wise leakage scores
The concept-wise leakage scores defined in (6) and (9) provide more granular information on the leakage encoded in each concept. In Figure 17 we display the $CTL_{i}$ and $ICL_{i}$ scores for the pairs of models analysed in Figure 4 and Table 1. Note in particular that i) leakage is not necessarily learnt homogeneously across concepts, and ii) comparing two models with different levels of interpretability, certain concepts may exhibit comparable amounts of leakage (such as $c_{4}$ in the dSprites(0) example), while for others leakage may be significantly different. Concept-wise scores are sensitive to both phenomena, making them valuable indicators of per-concept risk upon intervention as well as important measures for model design.
@@
Appendix D Details on the correlation between leakage scores and intervention performance
The models considered for this computation are
@@ Appendix D Details on the correlation between leakage scores and intervention performance
Each class of models were trained over 5 folds, amounting to a total of 60 evaluated models for each of the former four datasets, and to a total of 70 models for each of the latter two datasets.
As an example, in Figure 18 we display the distributions of the Pearson $r$ and $p$-values between intervention performance and the CTL, ICL, OIS scores resulting from MC-sampling for the TabularToy(0.25) dataset.
Figure 18: Distributions of Pearson $r$ and corresponding $p$-values between the intervention score $S_{(int)}$ and the CTL, ICL and OIS leakage metrics for TabularToy(0.25).
@@ Appendix E Behaviour of CBMs at high $\lambda$
In soft and logit CBMs leakage is minimal for values of $\lambda>\lambda_{min}$, where $\lambda_{min}$ is specific of the considered model class (Figure 19). As one further increases $\lambda$, leakage scores remain essentially constant. At significantly higher $\lambda\gtrsim\lambda_{max}$ concept supervision may however become too strong, at the expense of task learning. In such cases,
@@
    leakage may raise again well above the minimal value. The soft models trained on 3dshapes(0) are a compelling example of this.
Finally, note that the minimal attainable leakage in logit CBMs is higher than in soft CBMs on a given dataset.
@@ Appendix F Performance of models with incomplete concept sets and misspecified head
In Figure 20 we report the task and concept accuracies of the soft CBMs trained on complete and incomplete sets of concepts discussed in Section 7. Note in particular the decrease in task accuracy at high $\lambda$ that manifests in case of incomplete concept sets. At low $\lambda$, leakage is sufficient to ensure task accuracy remains essentially the same as in the case of complete concept sets. Figure 21 presents the task and concept accuracies of the well-specified and misspecified soft CBMs evaluated in Figure 8 and discussed in Section 7.
@@
Figure 20: Concepts and task accuracies evaluated for the same models analysed in Figure 7 on datasets with complete and incomplete sets of concepts.
@@
Figure 21: Concepts and task accuracies evaluated for the same models with a linear head analysed in Figure 8 trained on datasets with linear and non-linear tasks as functions of concepts.
Appendix G Dimensional-dependent bias in MI estimators
MI estimators are affected by significant biases in high dimensions as a general consequence of the curse of dimensionality (Lord et al., 2018; Carrara and Ernst, 2019; Czyż et al., 2023; Gowri et al., 2024). This results in MI estimates that cannot be compared across dimensions, as such biases are dimension-dependent. To illustrate this issue for the KSG estimator we adopt in this work, we consider normally-distributed higher-dimensional variables with non-trivial correlations, for which MIs and entropies can be computed in closed form. To see the effect of the dimension on the normalised interconcept MI, we sample two $d-$dimensional random variables $(X,Y)\sim\mathcal{N}(0,\Sigma^{(IC)}(\rho))$, where the covariance takes the form, $\Sigma^{(IC)}=\begin{pmatrix}\text{I}_{d}&\rho\,\text{I}_{d}\\ \rho\,\text{I}_{d}&\text{I}_{d}\end{pmatrix},$ where $\text{I}_{d}$ denotes the identity matrix of dimension $d$, and $\rho\in(-1,1)$ measures the linear correlation between each component of $X$ and $Y$. To model concepts-task MI, we consider two variables $X\in\mathbb{R}^{d}$ and $Y\in\mathbb{R}$ drawn from a normal distribution with zero mean and covariance matrix
$\Sigma^{(CT)}=\begin{pmatrix}1&\cdots&0&\rho\\ \vdots&\ddots&\vdots&\vdots\\ 0&\cdots&1&\rho\\ \rho&\cdots&\rho&1\end{pmatrix},$ (16)
where all the components of $X$ are equally correlated with $Y$. Note that $\Sigma^{(CT)}$ is positive-definite only for $0\leq\rho<1/\sqrt{d}$, hence we will consider covariances only within this range.
For normally-distributed variables, entropy and MI can be computed in closed form (Cover and Thomas, 2005). Considering two variables $(X,Y)\sim\mathcal{N}(\mu,\Sigma)$ of dimensions $d_{X}$ and $d_{Y}$ respectively, and $\Sigma^{(IC)}=\begin{pmatrix}\Sigma_{X}&\widetilde{\Sigma}\\ \widetilde{\Sigma}&\Sigma_{Y}\end{pmatrix},$ the entropy and MI read,
$H(X)=\frac{d_{X}}{2}\left(1+\log 2\pi\right),\qquad\quad I(X,Y)=\frac{1}{2}\log\frac{|\Sigma_{X}|\,|\Sigma_{Y}|}{|\Sigma|}\,,$ (17)
where ||$\cdot$|| ⋅ | denotes the determinant. For interconcept MI with $d_{X}=d_{Y}=d$, this entails the following theoretic values for MI and normalised MI,
$I^{(IC)}(X,Y)=-\frac{d}{2}\log\left(1-\rho^{2}\right),\qquad\widetilde{I}^{(IC)}(X,Y)=\frac{I^{(IC)}(X,Y)}{\sqrt{H(X)H(Y)}}=-\frac{\log\left(1-\rho^{2}\right)}{1+\log 2\pi}\,,$ (18)
while for concepts-task MI with $d_{X}=d$ and $d_{Y}=1$,
$I^{(CT)}(X,Y)=-\frac{1}{2}\log\left(1-d\rho^{2}\right),\qquad\widetilde{I}^{(CT)}(X,Y)=\frac{I^{(CT)}(X,Y)}{H(Y)}=-\frac{\log\left(1-d\rho^{2}\right)}{1+\log 2\pi}\,.$ (19)
We compare these predicted behaviours with the MIs and normalised MIs obtained with the KSG estimator (Kraskov et al., 2004) on a range of correlations and dimensions. The results are displayed in Figure 22, where the theoretic behaviours for different ground-truth correlations are represented with dashed lines.
Although monotonically related to the theoretic values, the estimated normalised MIs are pushed towards higher and higher values as one increases the dimension, regardless of the ground-truth correlation. This prevents comparing quantities estimated on concept representations of different dimensions as required by the definitions of the CTL and ICL scores in (6)-(8), while still enabling one to compare MIs and normalised MIs for concept representations of the same dimension. Note also that these results cannot be used as calibration curves to account for such biases, since these Gaussian models only incorporate linear correlations, while leakage more generally manifests as shared information between two variables.
Figure 22: Results for the Gaussian experiment illustrating the bias in estimates of MI and normalised MI depending on the dimension and covariance.
OIS is affected by the same type of biases.
As discussed in Section 4, the OIS is meant to be a rough estimate of interconcept leakage, and hence of interconcept MI when ground-truth interconcept correlations are negligible. In Figure 23 we display the OIS scores for the 16-dimensional concept vector representations $\bm{\hat{c}^{w}}_{i}$ in a number of CEMs trained with different $\lambda$ and $p_{int}$. We note OIS is generally very high and affected by the same biases as the standard MI estimators, casting further doubts on its utility for higher-dimensional concept representations.
Figure 23: The OIS values computed on the vectors $\bm{\hat{c}^{w}}_{i}$ learnt by CEMs with low and high $\lambda$ and $p_{int}$, on TabularToy(0.25), dSprites(0) and 3dshapes(0).
Appendix H Further results on CEM non-interpretability
Embedding structure at non-vanishing $p_{int}$.
In Figures 25 and 26 we display the PCA projections of the weighted vectors $\bm{\hat{c}^{w}}_{1}$ with colouring based on the ground-truth value of the task label and of concepts 2 and 3 respectively, at non-vanishing $p_{int}$ and for low and high $\lambda$.
Figure 24: Alignment leakage score for CEMs trained on the TabularToy(0.25) dataset with low and high $\lambda$ and $p_{int}$.
As demonstrated in Section 8, CEMs generally encode high amounts of concepts-task leakage, and from this perspective, concept accuracy is only a measure of how predictive the vectors $\bm{\hat{c}^{w}}_{i}$ are of concept $i$, fully unable to capture the additional information about the task that the embeddings encode. Furthermore, as shown in Figure 12, interconcept leakage increases as one raises concept supervision. Thus $c_{acc}$ typically correlates with interconcept leakage in CEMs.
Intervention performance is not an indicator of leakage either – it is very high in models with severe leakage (Figure 27). In particular, CEMs where high concepts-task leakage is present in both $\bm{\hat{c}^{+}}_{i}$ and $\bm{\hat{c}^{-}}_{i}$ are able to achieve a vanishing $S_{(int)}$. This motivates the definition and adoption of more sensitive information-theoretic measures for CEMs, such as (10), (11) and (12).
@@ Appendix H Further results on CEM non-interpretability
Figure 27: The $S_{(int)}$ score for CEMs with low and high $\lambda$ and $p_{int}$.
Appendix I Alignment leakage in CEMs
@@
A subtle leakage effect may in principle arise when exposing a model like CEMs to interventions during training. For each data point $n\in\{1,\dots,N\}$ with ground-truth value $c_{i}^{(n)}$ for concept $i$, we denote by aligned (unaligned) vector the one between $\hat{c}^{+(n)}_{i}$ or $\hat{c}^{-(n)}_{i}$ which is aligned (unaligned) with the value of $c_{i}^{(n)}$: if $c_{i}^{(n)}=1$, $\hat{c}^{+(n)}_{i}$ ($\hat{c}^{-(n)}_{i}$) is the aligned (unaligned) vector, and vice versa $\hat{c}^{-(n)}_{i}$ ($\hat{c}^{+(n)}_{i}$) is the aligned (unaligned) vector for $c_{i}^{(n)}=0$. The meaning of this distinction is that after intervening on the $i$-th concept at training or test time, the weighted vector $\hat{c}^{w(n)}_{i}$ is equal to the aligned vector. A model exposed to interventions may thus learn to improve performance upon interventions by generating aligned vectors that are more predictive of the task than the unaligned ones. This form of leakage results in embeddings that have inhomogeneous reliability depending on the alignment with ground-truth concepts, leading to an overall reduction in model interpretability.
@@
This effect, which we label alignment leakage, is a specific manifestation of concepts-task leakage. To quantify it, we define the following information-theoretic score, which indicates to what extent aligned vectors are more predictive of the task than unaligned vectors across both positive and negative embeddings, as an excess in concepts-task normalised MIs as defined in (10),
@@
where $\hat{c}^{\pm\text{(aligned)}}$ denote the set of positive/negative vectors aligned with the corresponding ground-truth concept, and analogously for $\hat{c}^{\pm\text{(unaligned)}}$. $\widetilde{I}^{(\text{align})}$ takes values in (-2,2) in units of normalised MI; positive values indicate that the aligned vectors are more predictive of the task than the unaligned ones, while a value of zero corresponds to no alignment leakage.
We assess this form of leakage in CEMs trained at different values of $\lambda$ and $p_{int}$. While in models trained on dSprites(0) and 3dshapes(0) it does not appear, we find $\widetilde{I}^{(\text{align})}$ scores incompatible with zero for models trained on TabularToy(0.25) (Figure 24). Alignment leakage is stronger in models at high $\lambda$ and $p_{int}$, supporting the intuition that such a detrimental effect correlates with exposure to interventions during training.

### v2

@@ A.1 Datasets
For additional details on these datasets, we refer the reader to Espinosa Zarlenga et al. (2023a).
@@
This dataset is constructed by sampling a 3-dimensional latent variable $z\sim\mathcal{N}(0,\Sigma(\delta))$, with correlation matrix $\Sigma(\delta)_{ij}=\delta_{ij}+\delta(1-\delta_{ij})$ with $i,j=1,2,3$, where $\delta_{ij}$ is the Kronecker symbol and $\delta\in(-1,1)$ indicating the correlation between components. The 7-dimensional input $x$ is a trigonometric function of the latent variables (see Mahinpei et al., 2021), while the three binary concepts are defined as the sign of the latent variables, $c_{i}=(z_{i}>0)$. The binary task label is the following linear function of concepts,
@@
For our experiments we generate a dataset of size 10,000, with a 0.7/0.2/0.1 train / validation / test ratio. For Figures 2 and 3 we consider a simplified version of TabularToy($\delta$) with only the first two concepts $c_{1}$ and $c_{2}$ for each input, and binary task ${y=(c_{1}+c_{2}\geq 1)}$.
@@
In TabularToy(0.25), we removed $c_{3}$ from the three concepts. In dSprites(0), we removed $c_{4}$ and $c_{5}$ from the five concepts. In 3dshapes(0), we removed $c_{3}$ and $c_{6}$ from the six concepts.
@@
We provide here more details on the experiments to assess the effects of a misspecified final head discussed in Section 7. To generate non-linear concepts-task dependences in TabularToy($\delta$), we added a non-linear term to the original linear task in (13), resulting in the binary classification problem,
@@
The original tasks in (14) and (15) for dSprites($\gamma$) and 3dshapes($\gamma$) are non-linear, however a linear classifier is able to reach essentially perfect task accuracy (Table 3). To induce a sizeable misspecification, we thus added further higher-order non-linear terms. The tasks implemented for our experiments on dSprites(0) and 3dshapes(0) are:
@@ A.2 Model architectures and training
On TabularToy($\delta$) we used as concept encoder a 4-layer leaky-ReLU MLP with activations $\{7,64,64,3\}$, and a linear classifier head. On dSprites($\gamma$) and 3dshapes($\gamma$) we used a ResNet-18 encoder, and a 4-layer ReLU MLP with activations $\{k,64,64,\ell\}$. We trained CEMs with 16-dimensional embeddings.
We trained all models for 200 epochs, using the Adam optimiser with learning rate $10^{-3}$ and momentum 0.9. For hard models, we trained the encoder and the classifier head for 200 and 20 epochs respectively. We trained with batch sizes 512 for TabularToy($\delta$) and 32 for dSprites($\gamma$) and 3dshapes($\gamma$).
A.3 Score evaluations
Except for Figure 14, all the displayed scores are evaluated on test sets. The reported intervention performances were obtained using a random policy: for each data point, the next concept to intervene on was selected uniformly at random from the set of concepts that had not yet been intervened on. In Figures 3, 4 and 16 we show results for single models, and the displayed means and 95% confidence intervals refer to a 5-fold evaluation of each score, including intervention performance. In the rest of the paper, we repeated the evaluation of each score 5 times for individual models, and we display means and 95% confidence intervals representing the distribution of the mean upon 5-fold training for each model class. MIs and entropies were estimated via the KSG estimator (Kraskov et al., 2004) with 3 nearest neighbours. The OIS and NIS were evaluated using their original implementations and default settings from Espinosa Zarlenga et al. (2023a).
Appendix B Undesired features of the Niche Impurity Score
@@
Because it is computed as an AUC, NIS = 1/2 indicates that concept representation subsets are not predictive of the value of concepts, while NIS = 1 indicates maximal higher-order interconcept leakage. For a fair comparison of our scores with NIS, we plotted $\widetilde{\text{NIS}}=2(\text{NIS}-1/2)$, so that scores of 0 and 1 correspond to no and maximal leakage respectively.
Evaluating NIS on the pairs of models considered in Figure 4, this score appears to be anticorrelated with intervention performance and leakage (Figure 16). Furthermore, it takes proportionally high values regardless of the model and dataset. This is confirmed by its behaviour on hard CBMs (Figure 15): the NIS is non-vanishing and high in all cases. These undesirable features indicate that NIS is not a suitable measure of leakage.
@@
Figure 16: $\widetilde{\text{NIS}}$ and leakage scores for the pairs of models assessed in Figure 4 and Table 1.
@@ Appendix C Concept-wise leakage scores
The concept-wise leakage scores defined in (6) and (9) provide more granular information on the leakage encoded in each concept. In Figure 17 we display the $CTL_{i}$ and $ICL_{i}$ scores for the pairs of models analysed in Figure 4 and Table 1. Note in particular that (i) leakage is not necessarily learnt homogeneously across concepts, and (ii) comparing two models with different levels of interpretability, certain concepts may exhibit comparable amounts of leakage (such as $c_{4}$ in the dSprites(0) example), while for others leakage may be significantly different. Concept-wise scores are sensitive to both phenomena, making them valuable indicators of per-concept leakage risk upon intervention and therefore useful measures for model design.
@@
Appendix D Details of the correlation between leakage scores and intervention performance
The models considered for this computation were
@@ Appendix D Details of the correlation between leakage scores and intervention performance
Each class of models was trained over 5 folds, amounting to a total of 60 evaluated models for each of the first set of four datasets, and to a total of 70 models for each of the second set of two datasets. We use Rubin’s rule (see Rubin, 2004; Barnard and Rubin, 1999) to combine within-imputation and between-imputation variances and obtain the best-estimate Pearson $r$ and $p$-value.
@@ Appendix E Behaviour of CBMs at high $\lambda$
In soft and logit CBMs leakage is minimal for values of $\lambda>\lambda_{min}$, where $\lambda_{min}$ is specific to the considered model class (Figure 18). As $\lambda$ is increased leakage scores remain essentially constant to start. However, at higher values of $\lambda\gtrsim\lambda_{max}$ concept supervision may become too strong, at the expense of task learning. In such cases,
@@
    leakage may raise again, to well above the minimal value. The soft models trained on 3dshapes(0) in Figure 18 are a compelling example of this.
Finally, note that the minimal attainable leakage in logit CBMs is higher than in soft CBMs for any given dataset and choice of $\lambda$, in line with the expectations.
@@ Appendix F Performance of models with incomplete concept sets and misspecified head
In Figure 19 we show the task and concept accuracies of the soft CBMs trained on the complete and incomplete sets of concepts discussed in Section 7. Note in particular the decrease in task accuracy at high values of $\lambda$ that occurs in the case of incomplete concept sets. At low $\lambda$, leakage is sufficient to ensure task accuracy remains essentially constant and approximately equivalent to that of the case of complete concept sets. Figure 20 shows the task and concept accuracies of the well-specified and misspecified soft CBMs evaluated in Figure 8 and discussed in Section 7.
@@
Figure 19: Concepts and task accuracies for the models analysed in Figure 7 on datasets with complete and incomplete sets of concepts.
@@
Figure 20: Concepts and task accuracies evaluated for models with a linear head analysed in Figure 8 trained on datasets with linear and non-linear tasks as functions of concepts.
@@
Appendix G Further results on CEM non-interpretability
Embedding structure at non-vanishing $p_{int}$.
In Figures 21 and 22 we display the PCA projections of the weighted vectors $\bm{\hat{c}^{w}}_{1}$ with colouring based on the ground-truth value of the task label and of concepts 2 and 3 respectively, at non-vanishing values of $p_{int}$ and for low and high values of $\lambda$.
Supportive evidence of leakage in CEMs.
The general dependence of CEMs on leaked information about the task is further substantiated by the fact that task accuracy is insensitive to the number of used concepts, as demonstrated in the experiments reported in Espinosa Zarlenga et al. (2022), Appendix 8. When training CEMs on the CUB dataset with up to 90% of the annotated concepts removed, the task accuracy remains essentially the same as that of models trained on all the concepts. Along the same lines, the rigorous analysis carried out in Espinosa Zarlenga et al. (2022) using the information bottleneck paradigm evidences that in CEMs there is no information bottleneck and that information can freely flow from the input to the task prediction. This result stands as a fundamental proof that the high-dimensional concept representations $\bm{\hat{c}^{w}}_{i}$ generally encode an arbitrary amount of information about the task, meeting the only requirement of being predictive of the real-valued concepts $c_{i}$.
Regarding interconcept leakage, concept embeddings were found to be very predictive of the value of the other concepts also in Espinosa Zarlenga et al. (2022), Appendix 9.
No functional $y(c)$ interpretability.
In CEMs, the final head takes as input the concatenated vectors $\bm{\hat{c}^{w}}_{i}$. This limits the interpretability of the learnt functional dependence $y=f(c)$, even when using an interpretable head such as a linear layer. Since the final head reasons at the level of individual vector components (which themselves are not interpretable), the only possible statements are of the form: ”the $m$-th component of the $i$-th concept vector is more relevant for the task prediction than the $n$-th component of the $j$-th concept vector.”
As demonstrated in Section 8, CEMs generally encode high amounts of concepts-task leakage, and from this perspective, concept accuracy is only a measure of how predictive the vectors $\bm{\hat{c}^{w}}_{i}$ are of concept $i$, regardless of the leaked information about the task in the weighted embeddings. Furthermore, as shown in Figure 12, interconcept leakage increases with concept supervision. Thus $c_{acc}$ typically correlates with interconcept leakage in CEMs.
Intervention performance is not an indicator of leakage either; indeed it can be high in models with severe leakage (see Figure 23). In particular, CEMs in which high concepts-task leakage is present in both $\bm{\hat{c}^{+}}_{i}$ and $\bm{\hat{c}^{-}}_{i}$ are able to achieve a vanishing $S_{int}$. This motivates the definition and adoption of more sensitive information-theoretic measures for CEMs, such as (10), (11) and (12).
Figure 23: The $S_{int}$ score for CEMs with low and high values of $\lambda$ and $p_{int}$.
Appendix H Alignment leakage in CEMs
@@
Figure 24: Alignment leakage score for CEMs trained on the TabularToy(0.25) dataset with low and high values of $\lambda$ and $p_{int}$.
@@
A subtle leakage effect may in principle arise when exposing a model like CEMs to interventions during training. For each data point $n\in\{1,\dots,N\}$ with ground-truth value $c_{i}^{(n)}$ for concept $i$, we denote by aligned (unaligned) vector the one between $\hat{c}^{+(n)}_{i}$ and $\hat{c}^{-(n)}_{i}$ which is aligned (unaligned) with the ground-truth value of $c_{i}^{(n)}$: if $c_{i}^{(n)}=1$, $\hat{c}^{+(n)}_{i}$ ($\hat{c}^{-(n)}_{i}$) is the aligned (unaligned) vector, and vice versa $\hat{c}^{-(n)}_{i}$ ($\hat{c}^{+(n)}_{i}$) is the aligned (unaligned) vector if $c_{i}^{(n)}=0$. The meaning of this distinction is that after intervening on the $i$-th concept at training or test time, the weighted vector $\hat{c}^{w(n)}_{i}$ is equal to the aligned vector. A model exposed to interventions may thus learn to improve performance upon interventions by generating aligned vectors that are more predictive of the task than the unaligned ones. This form of leakage results in embeddings that have inhomogeneous reliability depending on the alignment with ground-truth concepts, leading to an overall reduction in model interpretability.
@@
This effect, which we label alignment leakage, is a specific manifestation of concepts-task leakage. To quantify it, we define the following information-theoretic score, which indicates the extent to which aligned vectors are more predictive of the task than unaligned vectors across both positive and negative embeddings, as an excess in the concepts-task normalised MIs defined in (10). Specifically, we define:
@@
where $\hat{c}^{\pm\text{(aligned)}}$ denote the set of positive/negative vectors aligned with the corresponding ground-truth concept, and analogously for $\hat{c}^{\pm\text{(unaligned)}}$. $\widetilde{I}^{(\text{align})}$ takes values in (-2,2) in units of normalised MI. Positive values indicate that the aligned vectors are more predictive of the task than the unaligned ones, while a value of zero corresponds to no alignment leakage.
We assessed this form of leakage in CEMs trained at different values of $\lambda$ and $p_{int}$. While in models trained on dSprites(0) and 3dshapes(0) it does not appear, we find significantly non-zero scores for models trained on TabularToy(0.25) (Figure 24). Alignment leakage is stronger in models at high values of $\lambda$ and $p_{int}$, supporting the intuition that this detrimental effect is associated with exposure to interventions during training.